# Generalized Guarantees for Variational Inference in the Presence of Even and Elliptical Symmetry

### v1

@@ Abstract
We extend several recent results providing symmetry-based guarantees for variational inference (VI) with location-scale families. VI approximates a target density $p$ by the best match $q^{*}$ in a family $\mathcal{Q}$ of tractable distributions that in general does not contain $p$. It is known that VI can recover key properties of $p$, such as its mean and correlation matrix, when $p$ and $\mathcal{Q}$ exhibit certain symmetries and $q^{*}$ is found by minimizing the reverse Kullback-Leibler divergence. We extend these guarantees in two important directions. First, we provide symmetry-based guarantees for a broader family of divergences, highlighting the properties of variational objectives under which VI provably recovers the mean and correlation matrix. Second, we obtain further guarantees for VI when the target density $p$ exhibits even and elliptical symmetries in some but not all of its coordinates. These partial symmetries arise naturally in Bayesian hierarchical models, where the prior induces a challenging geometry but still possesses axes of symmetry. We illustrate these theoretical results in a number of experimental settings.
@@ 1 INTRODUCTION
Variational inference (VI) is a popular methodology for Bayesian inference and probabilistic machine learning (Jordan et al.,, 1999; Wainwright and Jordan,, 2008; Blei et al.,, 2017). The modus operandi of VI is to posit a family of tractable distributions $\mathcal{Q}$ and find within this family the best approximation to a target distribution $p$. VI is typically presented as a scalable alternative to more classical algorithms, such as Markov chain Monte Carlo (MCMC, Robert and Casella,, 2004). While VI can quickly obtain an approximation within a constrained family $\mathcal{Q}$, it is often unclear how well this solution approximates $p$ (e.g., Yao et al.,, 2018; Giordano et al.,, 2018; Talts et al.,, 2018; Huggins et al.,, 2020; Dhaka et al.,, 2021). For this reason it is useful to understand the conditions under which VI returns approximations that are provably accurate. This paper contributes to a growing body of work on this subject (e.g., Wang and Blei,, 2018; Katsevich and Rigollet,, 2024; Margossian and Saul,, 2025).
Most approaches to VI seek an approximation in $\mathcal{Q}$ by minimizing the reverse Kullback-Leibler (KL) divergence. But VI can be formulated with other divergences, some of which possess attractive properties, such as the forward KL-divergence (Naesseth et al.,, 2020; Vehtari et al.,, 2020) and the Rényi divergences (Li and Turner,, 2016; Dieng et al.,, 2017; Daudel et al.,, 2023). It is known that different divergences, when minmized, can return different solutions, even when the optimization is carried over the same family $\mathcal{Q}$. For this reason, it is of interest to understand which positive guarantees for VI hold across a range of divergences and are not sensitive to a particular choice of objective function. This paper considers this question for a family of divergences closely related to $f$-divergences (Rényi,, 1961).
Our work extends recent guarantees when $\mathcal{Q}$ is a family of location-scale distributions and the best variational approximation is found by minimizing the reverse KL divergence. Here it is known, under certain conditions, that VI recovers both the mean and correlation matrix of $p$ whenever $p$ is, respectively, even and elliptically symmetric (Margossian and Saul,, 2025). We generalize these results in two directions. First, we obtain similar guarantees for VI with a broader class of divergences. Second, we show that if $p$ exhibits symmetries in some but not all of its coordinates, then VI recovers the partial mean and correlations along these coordinates. Of special note is that such partial symmetries arise in models with hierarchical priors.
@@
Our work most closely builds on recent guarantees for VI in the presence of symmetries, despite misspecifications in the family $\mathcal{Q}$ (Margossian and Saul,, 2025). Earlier studies have also demonstrated VI’s ability to recover the mean, both empirically (e.g., MacKay,, 2003; Giordano et al.,, 2018) and theoretically, in pre-asymptotic regimes (e.g., Katsevich and Rigollet,, 2024), and others have obtained positive results when VI is used to maximize a marginal likelihood (Jordan et al.,, 1999; Li and Turner,, 2016) or construct frequentist estimators (Wang and Blei,, 2018; Alquier and Ridgway,, 2020; Yang et al.,, 2020; Zhang and Gao,, 2020). On the other hand, many studies have proven negative results for VI, particularly when it is used to quantify uncertainty in $p$ (MacKay,, 2003; Turner and Sahani,, 2011; Giordano et al.,, 2018; Margossian and Saul,, 2023; Margossian et al.,, 2025). A complementary line of work examines post-hoc diagnostics to assess the quality of VI—for example using importance sampling (Yao et al.,, 2018; Vehtari et al.,, 2024) or error bounds based on the Wasserstein distance (Huggins et al.,, 2020; Biswas and Mackey,, 2023).
Previous work has also explored alternative objective functions for VI. Several studies have shown that when different divergences are minimized, the approximations from VI yield different quantifications of uncertainty and different estimators of the marginal likelihood (Li and Turner,, 2016; Daudel et al.,, 2023; Margossian et al.,, 2025). Our results provide a counterpoint of sorts: we show that in the presence of certain symmetries, it may not matter which divergence is minimized, and that variational approximations from different divergences may all recover the mean and the correlation matrix of $p$, even when other properties of $p$ are poorly estimated.
Finally, our study of partial symmetries relates to a large literature on the geometry of posteriors in hierarchical models and the interplay of this geometry with inference algorithms (e.g. Neal,, 2001; Papaspiliopoulos et al.,, 2007; Betancourt and Girolami,, 2015).
@@ 2.1 Objective functions for VI
Divergence Notation $\varphi(t)$ Decreasing Linear
(Reverse) Kullback-Leibler $\text{KL}(q||p)$ t-t
[…]
(Forward) Kullback-Leibler $\text{KL}(p||q)$ $e^{t}t$
Squared Hellinger distance $H(p||q)$ $\left(e^{\frac{1}{2}t}-1\right)^{2}$
Total variation distance $\text{TV}(p,q)$ $\left|e^{t}-1\right|$
Table 1: Examples of common $\varphi$-divergences. Our theoretical results require that $\varphi$ is convex and differentiable. We obtain further guarantees when $\varphi$ is decreasing, and even further ones when $\varphi$ is linear.
@@
There exist several alternatives to the reverse KL-divergence. Certain algorithms approximately minimize the forward KL-divergence (Naesseth et al.,, 2020; Vehtari et al.,, 2020). A generalization of the KL-divergences is provided by the $\alpha$-divergence (Li and Turner,, 2016; Dieng et al.,, 2017; Daudel et al.,, 2023), which interpolates between the reverse and forward KL-divergences, respectively, in the limits $\alpha\!\to\!0$ and $\alpha\!\to\!1$ when the $\alpha$-divergence is defined as in Cichocki and Amari, (2010). It is well known that different divergences yield different variational approximations when the family $\mathcal{Q}$ is restricted and $p\notin\mathcal{Q}$.
Other measures of discrepancy between distributions include the Hellinger distance from information theory and the total variation distance, which is used to study MCMC (e.g. Roberts and Rosenthal,, 2004). In practice, these objective functions are not used for VI because they are too difficult to compute in high dimensions. Even for idealized objective functions, however, it remains of theoretical interest to understand when they yield an approximation $q$ that recovers statistical properties of $p$.
@@
($\varphi$-divergence) We refer to a divergence as a $\varphi$-divergence if it can be written as
where (i) $\varphi:\mathbb{R}\to\mathbb{R}$, (ii) $\varphi(0)=0$, (iii) $\varphi$ is convex, and (iv) $\varphi$ is differentiable.
Table 1 provides examples of common divergences that satisfy this definition. The family of $\varphi$-divergences is closely related to the family of $f$-divergences (Rényi,, 1961). We work with $\varphi$-divergences in this paper because they include the $\alpha$-divergences $D_{\alpha}(p||q)$ for $\alpha\!\in\!(0,1)$ and because they lead to clearer proofs in the setting where $p$ is log concave. We examine the connection between $\varphi$-divergences and $f$-divergences further in Appendix A.
Our theoretical analysis will leverage certain key properties of $\varphi$-divergences. We require that the function $\varphi$ is differentiable, an assumption that is tacitly made when practitionners use stochastic optimization in VI. We obtain our strongest results when $\varphi$ is convex and, additionally, when it is monotone decreasing. When these properties hold, we are able to identify settings where the $\varphi$-divergence has a unique global minimizer with respect to the variational parameters of $q$. Finally, for one of our results on partial symmetries, we additionally assume that $\varphi$ is linear; of all the divergences in Table 1, only the reverse KL divergence has this property. Table 1 summarizes, for each divergence, which of these additional regularity conditions are satisfied.
@@ 2.2 Even, elliptical, and partial symmetries
where $C$ is a correlation matrix. This is an extension of the well-known funnel (Neal,, 2001), in which $C$ is diagonal. The geometry of the funnel is typical of hierarchical priors and prone to frustrate many inference algorithms. But the funnel also has certain partial symmetries: the conditional distribution $p(\theta|\tau)$ is elliptically symmetric with a point of even symmetry and a normalized covariance matrix (proportional to $C$) that do not depend on the variable $\tau$. Later we will show that VI provably recovers the mean and the correlations along $\theta$ when $\mathcal{Q}$ is a family of elliptical distributions—even when VI misestimates the mean for $\tau$ (Figure 1).
@@ 2.3 Location-scale families
Location-scale families are prominent in statistics and probabilistic modeling. Examples include the Gaussian, Laplace, Student-t, and Cauchy distributions. They are also popular choices for VI (e.g., Ranganath et al.,, 2014; Kingma and Welling,, 2014; Kucukelbir et al.,, 2017; Cai et al., 2024b, ), where $\mathcal{Q}$ is often taken to be the family of Gaussians. Throughout the paper, we assume that the base distribution $q_{0}$ is spherically symmetric about the origin (and therefore also even symmetric about the origin).
@@ 2.4 Regularity conditions on $p$ and $q$
Our theoretical analysis is aided by imposing certain regularity conditions on $p$ and $q$. First, for all of our results, we require that $p$ is differentiable, and more generally, that for each $q\!\in\!\mathcal{Q}$ we can interchange the order of differentiation and integration when differentiating $D_{\varphi}(p||q)$. The formal conditions for this assumption are provided by the dominated convergence theorem (e.g., Billingsley,, 1995). This assumption is needed to minimize $D_{\varphi}(p||q)$ via stochastic optimization, and in particular to compute Monte Carlo estimates of gradients with respect to the variational parameters.
@@ 3.1 Guarantees from even symmetry
First we generalize earlier guarantees for VI with even symmetries (Margossian and Saul,, 2025) to the broader family of $\varphi$-divergences.
@@
Let $\mathcal{Q}$ be a location family and let $D_{\varphi}$ be a $\varphi$-divergence. If $p$ is even symmetric about $\mu$, then a stationary point of $D_{\varphi}(p||q_{\nu})$ occurs at $\nu\!=\!\mu$. Furthermore, if $\varphi$ is strictly decreasing, and $p$ somewhere-strictly log concave over $\mathbb{R}^{d}$, then $\nu\!=\!\mu$ is a unique minimizer of $D_{\varphi}(p||q_{\nu})$.
@@
Let $\zeta=z\!-\!\nu$. By definition, $q_{\nu}(z)=q_{0}(\zeta)$, so that we can write
@@
$=\int\frac{\nabla_{\nu}p(\nu+\zeta)}{q_{0}(\zeta)}\varphi^{\prime}\left(\log\frac{p(\nu+\zeta)}{q_{0}(\zeta)}\right)q_{0}(\zeta)\text{d}\zeta$
$=\int\frac{\nabla_{\zeta}p(\nu+\zeta)}{q_{0}(\zeta)}\varphi^{\prime}\left(\log\frac{p(\nu+\zeta)}{q_{0}(\zeta)}\right)q_{0}(\zeta)\text{d}\zeta,$ (9)
@@
Now suppose that $\varphi$ is strictly monotone decreasing and $p$ is somewhere-strictly log concave on $\mathbb{R}^{d}$. Then we show in Lemma 16 of Appendix B.1 that $D_{\varphi}(p||q_{\nu})$ is strictly convex in $\nu$, and hence the stationary point at $\mu\!=\!\nu$ is a unique minimizer. ∎
Theorem 10 allows for several misspecifications in the variational approximation. For example, $q$ may be factorized while $p$ is not, or $q$ and $p$ may behave differently in their tails. Fig. 2 (left) illustrates VI’s ability to recover the mean when $p(z)$ is a 2-dimensional student-t distribution with 5 degrees of freedom and correlation $\rho\!=\!0.7$, and $\mathcal{Q}$ is a family of factorized Gaussians. Minimizing the $\varphi$-divergences in Table 1 (via grid search), we find that they all yield the same, exact estimate of the mean $\nu$. Interestingly, the exact mean is recovered for all five divergences; this is in fact a stronger result than what the theorem guarantees. While the theorem states that the mean-matching solution is a stationary point for all divergences, only the first two divergences in Table 1—the reverse KL and the $\alpha$-divergences with $\alpha\in(0,1)$—satisfy the additional assumption that $\varphi$ is monotone decreasing and which guarantees that $\nu\!=\!\mu$ is a unique minimizer.

### v2

@@ Abstract
We extend several recent results providing symmetry-based guarantees for variational inference (VI) with location-scale families. VI approximates a target density $p$ by the best match $q^{*}$ in a family $\mathcal{Q}$ of tractable distributions that in general does not contain $p$. It is known that VI can recover key properties of $p$, such as its mean and correlation matrix, when $p$ and $\mathcal{Q}$ exhibit certain symmetries and $q^{*}$ is found by minimizing the reverse Kullback-Leibler divergence. We extend these guarantees in two important directions. First, we provide symmetry-based guarantees for $f$-divergences, a broad class that includes the reverse and forward Kullback-Leibler divergences and the $\alpha$-divergences. We highlight properties specific to the reverse Kullback-Leibler divergence under which we obtain our strongest guarantees. Second, we obtain further guarantees for VI when the target density $p$ exhibits even and elliptical symmetries in some but not all of its coordinates. These partial symmetries arise naturally in Bayesian hierarchical models, where the prior induces a challenging geometry but still possesses axes of symmetry. We illustrate these theoretical results in a number of experimental settings.
@@ 1 INTRODUCTION
Variational inference (VI) is a popular methodology for Bayesian inference and probabilistic machine learning (Jordan:1999; Wainwright:2008; Blei:2017). The modus operandi of VI is to posit a family of tractable distributions $\mathcal{Q}$ and find within this family the best approximation to a target distribution $p$. VI is typically presented as a scalable alternative to more classical algorithms, such as Markov chain Monte Carlo (MCMC, Robert:2004). While VI can quickly obtain an approximation within a constrained family $\mathcal{Q}$, it is often unclear how well this solution approximates $p$ (e.g., Yao:2018; Giordano:2018; Talts:2018; Huggins:2020; Dhaka:2021). For this reason it is useful to understand the conditions under which VI returns approximations that are provably accurate. This paper contributes to a growing body of work on this subject (e.g., Wang:2018; Katsevich:2024; Margossian:2025robust).
Most approaches to VI seek an approximation in $\mathcal{Q}$ by minimizing the reverse Kullback-Leibler (KL) divergence. But VI can be formulated with other divergences, some of which possess attractive properties, such as the forward KL-divergence (Wainwright:2008; Naesseth:2020; Vehtari:2020) and the $\alpha$-divergences (Li:2016; Dieng:2017; Daudel:2023). It is known that different divergences, when minmized, can return different solutions, even when the optimization is carried over the same family $\mathcal{Q}$. For this reason, it is of interest to understand which positive guarantees for VI hold across a range of divergences and are not sensitive to a particular choice of objective function. Here, we consider this question for $f$-divergences (Renyi:1961).
Our work extends recent guarantees when $\mathcal{Q}$ is a family of location-scale distributions and the best variational approximation is found by minimizing the reverse KL divergence. Here it is known, under certain conditions, that VI recovers both the mean and correlation matrix of $p$ whenever $p$ is, respectively, even and elliptically symmetric (Margossian:2025robust). We generalize these results in two directions. First, we obtain similar guarantees for VI with a broader class of divergences. Second, we show that if $p$ exhibits symmetries in some but not all of its coordinates, then VI recovers the partial mean and correlations along these coordinates. Of special note is that such partial symmetries arise in models with hierarchical priors.
@@
Much of the literature on VI focuses on minimizing the reverse KL-divergence. Here, our work most closely builds on recent guarantees for VI in the presence of symmetries, despite misspecifications in the family $\mathcal{Q}$ (Margossian:2025robust). Earlier studies have also demonstrated VI’s ability to recover the mean, empirically (e.g., Mackay:2003; Giordano:2018) and theoretically (e.g., Katsevich:2024), and others have obtained positive results when VI is used to maximize a marginal likelihood (Jordan:1999; Li:2016) or construct frequentist estimators (Wang:2018; Alquier:2020; Yang:2020; Zhang:2020). On the other hand, many studies have proven negative results for VI, particularly when it is used to quantify uncertainty in $p$ (Mackay:2003; Turner:2011; Giordano:2018; Margossian:2023; Margossian:2025uncertainty). A complementary line of work examines post-hoc diagnostics to assess the quality of VI—for example using importance sampling (Yao:2018; Vehtari:2024) or error bounds based on the Wasserstein distance (Huggins:2020; Biswas:2024).
Previous studies have also explored alternative objective functions for VI. For example, when the forward KL-divergence is minimized by VI, it is known that the optimal Gaussian approximation exactly recovers the target density’s mean and covariance. (Wainwright:2008). Studies have also shown that when different divergences are minimized, the approximations from VI yield different quantifications of uncertainty and different estimators of the marginal likelihood (Li:2016; Daudel:2023; Margossian:2025uncertainty). Our results provide a counterpoint of sorts: we show that in the presence of certain symmetries, variational approximations from different divergences may all recover the mean and the correlation matrix of $p$, even when other properties of $p$ are poorly estimated.
Finally, our study of partial symmetries relates to a large literature on the geometry of posteriors in hierarchical models and the interplay of this geometry with inference algorithms (e.g. Neal:2001; Papaspiliopoulos:2007; Betancourt:2015).
@@ 2.1 Objective functions for VI
Divergence Notation $f(u)$ $\varphi(v)$
(Reverse) Kullback-Leibler $\text{KL}(q||p)$ $\log u$ v-v
$\alpha$-divergence, $\alpha>0$ $\text{D}_{\alpha}(p||q)$ $\frac{u^{\alpha}-1}{\alpha(\alpha-1)}$ $\frac{e^{\alpha v}-1}{\alpha(\alpha-1)}$
(Forward) Kullback-Leibler $\text{KL}(p||q)$ $u\log u$ $ve^{v}$
Squared Hellinger distance $H(p||q)$ $(\sqrt{u}-1)^{2}$ $\left(e^{\frac{1}{2}v}-1\right)^{2}$
Total variation distance $\text{TV}(p,q)$ |u1||u-1| $\left|e^{v}-1\right|$
Table 1: Common $f$-divergences. The $f$-divergences can be reformulated with the $\varphi$ function when $v=\log u$. Our theoretical results requires $\varphi$ to be differentiable, which is verified by all the divergences in the table. We obtain stronger guarantees when $\varphi$ is convex and decreasing. In this table, these conditions are only verified by $\text{KL}(q||p)$. We further obtain results on partial symmetry for $\text{KL}(q||p)$ by exploiting the fact $\varphi$ is linear.
@@
There exist several alternatives to the reverse KL-divergence. Certain algorithms approximately minimize the forward KL-divergence (Naesseth:2020; Vehtari:2020). A generalization of the KL-divergences is provided by the $\alpha$-divergence (Li:2016; Dieng:2017; Daudel:2023), which interpolates between the reverse and forward KL-divergences, respectively, in the limits $\alpha\!\to\!0$ and $\alpha\!\to\!1$ when the $\alpha$-divergence is defined as in Cichocki:2010. Furthermore, for $\alpha=1/2$, minimizing the $\alpha$-divergence is equivalent to minimizing the squared Hellinger distance. Another useful measure of discrepancy between distributions is the total variation distance, used for MCMC (e.g. Roberts:2004). Many of the here discussed objective functions are not used for VI because they are too difficult to compute in high dimensions. Even for idealized objective functions, however, it remains of theoretical interest to understand when they yield an approximation $q$ that recovers statistical properties of $p$.
@@
($f$-divergence (Renyi:1961)) We refer to a divergence as an $f$-divergence if it can be written as
$D_{f}(p||q)=\int f\left(\frac{p(z)}{q(z)}\right)\text{d}z,$ (2)
where $f:\mathbb{R}^{+}\to\mathbb{R}$ is such that (i) $f$ is convex, (ii) $f(1)=0$ and (iii) $\text{lim}_{u\to 0^{+}}f(u)=f(0)$.
It will be useful to work on the log scale and reason about the properties of $\log p(z)$. To do so, we introduce the function $\varphi:\mathbb{R}\to\mathbb{R}$, defined as,
$\varphi\left(\log\frac{p(z)}{q(z)}\right):=f\left(\frac{p(z)}{q(z)}\right).$ (3)
We then write,
$D_{\varphi}(p||q):=\int\varphi\left(\log\frac{p(z)}{q(z)}\right)q(z)\text{d}z.$ (4)
For $\varphi$ defined as in eq. (3), $D_{\varphi}(p||q)=D_{f}(p||q)$. In this paper, we work with the formulation in eq. (4), noting that any results that apply to $D_{\varphi}(p||q)$ apply to the corresponding $D_{f}(p||q)$. Table 1 provides examples of common $f$-divergences with the corresponding $f$ and $\varphi$ functions.
Our theoretical analysis requires that the functions $f$ is differentiable, an assumption that is tacitly made when practitioners use stochastic optimization in VI. An immediate consequence of this assumption is that $\varphi$ is also differentiable. We obtain our strongest results when $\varphi$ is convex and monotone decreasing. When these properties hold, we are able to identify settings where the $f$-divergence has a unique global minimizer with respect to the variational parameters of $q$. These conditions on $\varphi$ are only verified by $\text{KL}(q||p)$ among the divergences in Table 1. Similarly, one of our results on partial symmetry requires $\varphi$ to be linear which is again only verified by $\text{KL}(q||p)$. These conditions, while sufficient, may not be necessary. Finding weaker conditions that apply more broadly to $f$-divergences is an open question.
@@ 2.2 Even, elliptical, and partial symmetries
where $C$ is a correlation matrix. This is an extension of the well-known funnel (Neal:2001), in which $C$ is diagonal. The geometry of the funnel is typical of hierarchical priors and prone to frustrate many inference algorithms. But the funnel also has certain partial symmetries: the conditional distribution $p(\theta|\tau)$ is elliptically symmetric with a point of even symmetry and a normalized covariance matrix (proportional to $C$) that do not depend on the variable $\tau$. Later we will show that VI provably recovers the mean and the correlations along $\theta$ when $\mathcal{Q}$ is a family of elliptical distributions—even when VI misestimates the mean for $\tau$ (Figure 1).
@@ 2.3 Location-scale families
Location-scale families are prominent in statistics and probabilistic modeling. Examples include the Gaussian, Laplace, Student-t, and Cauchy distributions. They are also popular choices for VI (e.g., Ranganath:2014; Kingma:2014; Kucukelbir:2017; Cai:2024), where $\mathcal{Q}$ is often taken to be the family of Gaussians. Throughout the paper, we assume that the base distribution $q_{0}$ is spherically symmetric about the origin (and therefore also even symmetric about the origin).
@@ 2.4 Regularity conditions on $p$ and $q$
Our theoretical analysis is aided by imposing certain regularity conditions on $p$ and $q$. First, for all of our results, we require that $p$ is differentiable, and more generally, that for each $q\!\in\!\mathcal{Q}$ we can interchange the order of differentiation and integration when differentiating $D_{\varphi}(p||q)$. The formal conditions for this assumption are provided by the dominated convergence theorem (e.g., Billingsley:1995). This assumption is needed to minimize $D_{\varphi}(p||q)$ via stochastic optimization, and in particular to compute Monte Carlo estimates of gradients with respect to the variational parameters.
@@ 3.1 Guarantees from even symmetry
First we generalize earlier guarantees for VI with even symmetries (Margossian:2025robust) to the broader family of $f$-divergences.
@@
Let $\mathcal{Q}$ be a location family and let $D_{f}$ be an $f$-divergence. If $p$ is even symmetric about $\mu$, then a stationary point of $D_{f}(p||q_{\nu})$ occurs at $\nu\!=\!\mu$. Furthermore, if $\varphi(v)=f\circ\exp(v)$ is convex and strictly decreasing, and $p$ somewhere-strictly log concave over $\mathbb{R}^{d}$, then $\nu\!=\!\mu$ is a unique minimizer of $D_{f}(p||q_{\nu})$.
@@
Let $\zeta=z\!-\!\nu$. By definition, $q_{\nu}(z)=q_{0}(\zeta)$. Using the formulation of the $f$-divergence in eq. (4), we have
@@
$=\int\frac{\nabla_{\nu}p(\nu+\zeta)}{p(\nu+\zeta)}\varphi^{\prime}\left(\log\frac{p(\nu+\zeta)}{q_{0}(\zeta)}\right)q_{0}(\zeta)\text{d}\zeta$
$=\int\frac{\nabla_{\zeta}p(\nu+\zeta)}{p(\nu+\zeta)}\varphi^{\prime}\left(\log\frac{p(\nu+\zeta)}{q_{0}(\zeta)}\right)q_{0}(\zeta)\text{d}\zeta,$ (11)
@@
Now suppose that $\varphi$ is convex, strictly monotone decreasing and $p$ is somewhere-strictly log concave on $\mathbb{R}^{d}$. Then we show in Lemma 15 of Appendix A.1 that $D_{\varphi}(p||q_{\nu})$ is strictly convex in $\nu$, and hence the stationary point at $\mu\!=\!\nu$ is a unique minimizer. ∎
Theorem 10 allows for several misspecifications in the variational approximation. For example, $q$ may be factorized while $p$ is not, or $q$ and $p$ may behave differently in their tails. Fig. 2 (left) illustrates VI’s ability to recover the mean when $p(z)$ is a 2-dimensional student-t distribution with 5 degrees of freedom and correlation $\rho\!=\!0.7$, and $\mathcal{Q}$ is a family of factorized Gaussians. Minimizing the $f$-divergences in Table 1 (via grid search), we find that they all yield the same, exact estimate of the mean $\nu$. Interestingly, the exact mean is recovered for all five divergences; this is in fact a stronger result than what the theorem guarantees. Indeed, the theorem states that the mean-matching solution is a stationary point. On the other hand, the conditions required for this stationary point to be a unique minimizer are not verified in this example because $p$ is not log concave (rather it is quasi concave) and $\varphi$ is not convex and decreasing for the divergences in Table 1, except for $\text{KL}(q||p)$.